PART I: PERT Consider the following table: ACTIVITY      PREC. ACT.              TIME ESTIMATES     (days)       &n

PART I: PERT

Consider the following table:

ACTIVITY      PREC. ACT.              TIME ESTIMATES     (days)        

to                       tml                 tp                       te

    A                      ---                           2                      3                      4                     

    B                      A                            2                      4                      6

    C                      A                            4                      6                      8

    D                      A                            4                      5                      6

    E                       B                            6                      7                      8

    F                       D                            1                      3                      5

    G                 E, C, F                        3                      5                      7

1.   Construct a PERT network. (20 points)

[NOTE: This is the most crucial part of the exercise. Make sure you have the correct network before proceeding to the next sections.]

2.   By using ES, EF, LF, LS, find the slack of each activity and clearly state the critical path. (10 points)

3.   Find the completion time ( ) and the standard deviation (σ) for the project— round off the standard deviation to the nearest hundredth: two decimal places. (20 points)

[NOTE: Sigma is only computed on the critical path.]

4.   What is the completion time at 68% and 95% confidence levels? (20 points)

5.   What is the likelihood of this project being completed in more than 20 days? (15 points)

6.   Between 16 and 18 days? (15 points)
Answer #1

Formula:

Expected Activity Duration using Beta Distribution / Formula = (O + (4 X M) + P)/6

Variance =Standard Deviation^2

Standard Deviation=(Pessimistic Time -Optimistic Time)/6

1 B C D E F G H I J
2 Activity Predecessor Optimistic time - O Most Likely Time - M Pessimistic Time - P Expected Activity Duration using Beta Distribution / Formula = (O + (4 X M) + P)/6 Variance Standard Deviation Activity on Critical Path - Yes or No
3 A --- 2 3 4 3 0.11 0.33 Yes
4 B A 2 4 6 4 0.44 0.67 Yes
5 C A 4 6 8 6 0.44 0.67 No
6 D A 4 5 6 5 0.11 0.33 No
7 E B 6 7 8 7 0.11 0.33 Yes
8 F D 1 3 5 3 0.44 0.67 No
9 G E,C,F 3 5 7 5 0.44 0.67 Yes
10 Total Variance of the Critical Path 1.11
11 Standard Deviation of the Critical Path 1.05

Formula:

1 B C D E F G H I J
2 Activity Predecessor Optimistic time - O Most Likely Time - M Pessimistic Time - P Expected Activity Duration using Beta Distribution / Formula = (O + (4 X M) + P)/6 Variance Standard Deviation Activity on Critical Path - Yes or No
3 A --- 2 3 4 =(D3+(4*E3)+F3)/6 =I3^2 =(F3-D3)/6 Yes
4 B A 2 4 6 =(D4+(4*E4)+F4)/6 =I4^2 =(F4-D4)/6 Yes
5 C A 4 6 8 =(D5+(4*E5)+F5)/6 =I5^2 =(F5-D5)/6 No
6 D A 4 5 6 =(D6+(4*E6)+F6)/6 =I6^2 =(F6-D6)/6 No
7 E B 6 7 8 =(D7+(4*E7)+F7)/6 =I7^2 =(F7-D7)/6 Yes
8 F D 1 3 5 =(D8+(4*E8)+F8)/6 =I8^2 =(F8-D8)/6 No
9 G E,C,F 3 5 7 =(D9+(4*E9)+F9)/6 =I9^2 =(F9-D9)/6 Yes
10 Total Variance of the Critical Path =SUM(H3+H4+H7+H9)
11 Standard Deviation of the Critical Path =SQRT(H10)

1)

Network diagram:

7 E 14 3 B 7 0 7 0 4 0 7 0 14 7 3 14 G 19 9 с 6 0 5 0 con w 5 0 14 14 A 3 19 0 3 0 3 0 3 D 8 8 F 11 3 5 3 3 3 11 3 14 6 11

2)

Legends are as stated below to find ES, EF, LS, LF and Total float of each activity

Activity Id Early Finish-Early Start + Duration Duration Early Start=Highest Early finish time of all predecessor activities

here, critical path is A-B-E-G as activities on this path have zero total float

3)

Project completion time=19 days

Standard Deviation =1.05 as shown in the excel sheet with formula

4)

Completion time at 95% confidence level

=(19-(1.96*1.05/Square root of 4)) to (19+(1.96*1.05/Square root of 4))

= (19-(1.96*1.05/2))to (19+(1.96*1.05/2))

= (19-1.03)to (19+1.03)

=17.97 to 20.03

Completion time at 68% confidence level

=(19-(0.468*1.05/Square root of 4)) to (19+(0.468*1.05/Square root of 4))

=(19-(0.468*1.05/2))to (19+(0.468*1.05/2))

= 18.754 to 19.245

5)

Project will take at least 19 days time to complete, so μ= 19 Days, Standard Deviation = 1.05

Given, X = 20 Days

Thus, Value of Z score = (X- μ)/ Standard Deviation = (20-19)/1.05 = 0.95 (Rounded up to 2 Decimal Numbers)

Thus, using Z table, The probability that the project will complete within 20 days = 82.894%

The probability that the project will take more than 20 days=(100-82.894)%=17.106%

6)

Project will take at least 19 days time to complete, so μ= 19 Days, Standard Deviation = 1.05

Given, X = 20 Days

Thus, Value of Z score = (X- μ)/ Standard Deviation = (20-19)/1.05 = 0.95 (Rounded up to 2 Decimal Numbers)

Thus, using Z table, The probability that the project will complete within 20 days = 82.894%

The probability that the project will take more than 20 days=(100-82.894)%=17.106%

Project will take at least 19 days time to complete, so μ= 19 Days, Standard Deviation = 1.05

Given, X = 16 Days

Thus, Value of Z score = (X- μ)/ Standard Deviation = (16-19)/1.05 = -2.86 (Rounded up to 2 Decimal Numbers)

Thus, using Z table, The probability that the project will complete within 16 days = 0.00212=.212%

Given, X = 18 Days

Thus, Value of Z score = (X- μ)/ Standard Deviation = (18-19)/1.05 = -0.95 (Rounded up to 2 Decimal Numbers)

Thus, using Z table, the probability that the project will complete within 18 days=.17106=17.106%

Thus, using Z table, the probability that the project will complete Between 16 and 18 days=(17.106-0.212)=16.89%